40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 26 17:50:22 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement 
    Reply  Wed Jan 27 18:19:45 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement 
       Reply  Thu Jan 28 00:29:30 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement PLL_OLG.pdfFreq_noise_comparison.pdf
          Reply  Mon Feb 1 18:11:25 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement Free_running_frequency_noise_comparison.pdf
       Reply  Fri Jan 29 14:43:47 2016, Steve, Update, Green Locking, Innolight laser is 10 years old inno1W.pdf
          Reply  Fri Jan 29 16:55:27 2016, rana, Update, Green Locking, Innolight laser is 10 years old 
             Reply  Sat Jan 30 00:12:22 2016, gautam, Update, Green Locking, Innolight laser is 10 years old RIN_comparison.pdf
                Reply  Sat Jan 30 09:56:24 2016, Koji, Update, Green Locking, Innolight laser is 10 years old 
                   Reply  Mon Feb 1 15:16:28 2016, gautam, Update, Green Locking, Innolight laser is 10 years old 
                      Reply  Mon Feb 1 15:43:18 2016, Koji, Update, Green Locking, Innolight laser is 10 years old 
                         Reply  Fri Feb 5 15:02:13 2016, gautam, Update, Green Locking, X-end NE cable 
          Reply  Fri Feb 12 11:10:49 2016, Steve, Update, Green Locking, Innolight laser is 10 years old inno1W.jpg
          Reply  Mon Mar 21 14:29:32 2016, Steve, Update, Calibration-Repair, 1W Innolight laser repair diagnoses 
             Reply  Thu Mar 24 07:56:09 2016, Steve, Update, Calibration-Repair, NO Noise Eater for 1W Innolight 
                Reply  Mon Apr 11 17:03:41 2016, Steve, Update, Calibration-Repair, 1W Innolight repair completed 
Message ID: 11963     Entry time: Sat Jan 30 00:12:22 2016     In reply to: 11962     Reply to this: 11964
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Innolight laser is 10 years old 

 

Quote:

I don't think there's any evidence that the noise eater is bad. That would change the behavior of the relaxation oscillation which is at 1 MHz ?

While I was investigating the AM/PM ratio of the Innolight, I found that there was a pronounced peak in the RIN at ~400kHz, which did not change despite toggling the noise eater switch on the front panel (see plot attached). The plot in the manual suggests the relaxation oscillations should be around 600kHz, but given that the laser power has dropped by a factor of ~3, I think it's reasonable that the relaxation oscillations are now at ~400kHz? 

Attachment 1: RIN_comparison.pdf  54 kB  | Show | Show all
ELOG V3.1.3-